# xLSTM-Mixer: Multivariate Time Series Forecasting by Mixing via Scalar Memories

### v2

@@ 6 Conclusion
Reproducibility Statement
All implementation details, including dataset descriptions, metric calculations, and experiment configurations, are provided in Sec. 4 and Sec. A.1. We make sure to exclusively use openly available software and datasets and provide the source code for full reproducibility at
https://github.com/mauricekraus/xLSTM-Mixer.
Acknowledgments
This work received funding from the EU project EXPLAIN, under the Federal Ministry of Education and Research (BMBF) (grant 01—S22030D). Furthermore, it was funded by the ACATIS Investment KVG mbH project “Temporal Machine Learning for Long-Term Value Investing” and the BMBF KompAKI project within the “The Future of Value Creation – Research on Production, Services and Work” program (funding number 02L19C150), managed by the Project Management Agency Karlsruhe (PTKA). The author of Eindhoven University of Technology received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. Furthermore, this work benefited from the HMWK project “The Third Wave of Artificial Intelligence - 3AI”.
Appendix A Appendix
A.1 Implementation Details
Experimental Details
@@ A.1 Implementation Details
Our codebase is implemented in Python 3.11, leveraging PyTorch 2.4  (Paszke et al., 2019) in combination with Lightning 2.4222https://lightning.ai/pytorch-lightning for model training and optimization. We used the custom CUDA implementation of Beck et al. (2024) for sLSTM [^3] https://github.com/NX-AI/xlstm which relies on the NVIDIA Compute Capability 8.0 or higher. Thus, our experiments were conducted on a single NVIDIA A100 80GB GPU. The majority of our baseline implementations, along with data loading and preprocessing steps, are adapted from the Time-Series-Library [^4] https://github.com/thuml/Time-Series-Library of Wang et al. (2024b). Additionally, for xLSTMTime we used code based on the official repository [^5] https://github.com/muslehal/xLSTMTime of Alharthi & Mahmood (2024).
Training and Hyperparameters
@@
We optimized xLSTM-Mixer for up to 60 epochs with a cosine-annealing scheduler with the Adam optimizer (Kingma & Ba, 2017), using $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and no weight decay. Hyperparameter tuning was conducted using Optuna (Akiba et al., 2019) with the choices provided in Tab. 4. We optimized for the L1 forecast error, also known as the Mean Absolute Error (MAE). To further stabilize the training process, gradient clipping with a maximum norm of $1.0$ was applied. All experiments were run with three different random seeds {2021, 2022, 2023}.
@@
Metrics
We follow common practice in the literature (Wu et al., 2021; Wang et al., 2024a) for maximum comparability and, therefore, evaluate long-term forecasting of all models on the mean absolute error (MAE), mean squared error (MSE), and for short-term forecasting, using the MAE, root mean squared error (RMSE), and mean absolute percentage error (MAPE). The metrics are averaged over all variates and computed as:
$\operatorname{MAE}(\bm{y},\bm{\hat{y}})$ $=\sum_{i=1}^{H}\left|y_{i}-\hat{y_{i}}\right|$ $\operatorname{MSE}(\bm{y},\bm{\hat{y}})$ $=\sum_{i=1}^{H}(y_{i}-\hat{y_{i}})^{2}$
$\operatorname{RMSE}(\bm{y},\bm{\hat{y}})$ $=\sqrt{\operatorname{MSE}(\bm{y},\bm{\hat{y}})}$ $\operatorname{MAPE}(\bm{y},\bm{\hat{y}})$ $=\frac{100}{H}\sum_{i=1}^{H}\frac{\left|y_{i}-\hat{y}_{i}\right|}{\left|y_{i}\right|+\epsilon},$
where $\bm{y}$ are the targets, $\bm{\hat{y}}$ the predictions, and $\epsilon$ a small constant added for numerical stability.
A.2 Outlook: Short-Term Time Series Forecasting
Having shown superior long-term forecasting accuracies in Sec. 4.1, we also provide an initial exploration of the effectiveness of xLSTM-Mixer to short-term forecasts. To this end, we compare it to applicable baselines on PEMS datasets with input lengths uniformly set to 96 and prediction lengths to 12. The results in Tab. 5 show that the performance of xLSTM-Mixer is competitive with existing methods. We provide the MAE, MAPE, and RMSE as is common practice.
Table 5: Short-term forecasting evaluation of xLSTM-Mixer and baselines on the multivariate PEMS datasets. A lower MAE, MAPE, or RMSE indicates a better prediction. The best result for each dataset is highlighted in bold red, while the second-best result is blue and underlined.
Models Recurrent MLP Transformer Convolutional
xLSTM-Mixer
@@ A.2 Outlook: Short-Term Time Series Forecasting
TimeMixer
DLinear
PatchTST
FEDFormer
Autoformer
MICN
TimesNet
1997a
PEMS03
MAE
15.71
16.59
18.65
14.63
19.70
18.95
19.00
18.08
15.71
16.41
MAPE
14.92
15.31
17.39
14.54
18.35
17.29
18.57
18.75
15.67
15.17
RMSE
24.82
26.47
31.73
23.28
32.35
30.15
30.05
27.82
24.55
26.72
PEMS08
MAE
16.56
17.44
20.34
15.22
20.26
20.35
20.56
20.47
17.76
19.01
MAPE
10.24
10.58
13.05
9.67
12.09
13.15
12.41
12.27
10.76
11.83
RMSE
26.65
28.13
31.90
24.26
32.38
31.04
32.97
31.52
27.26
30.65
  • a
    Configuration following Wu et al. (2021).
A.3 Full Results for Long-Term Forecasting
Tab. 6 shows the full results for long-term forecasting on all four separate forecast horizons.
Table 6: Full long-term forecasting results for Tab. 2. Avg is averaged from all four prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted in bold red, while the second-best result is blue and underlined. Wins for each model are shown at the bottom.
Models Recurrent MLP Transformer Convolutional
xLSTM-Mixer
xLSTMTime
LSTM
TimeMixer

### v3

@@ Appendix A Societal Impact
Appendix B Rationale for Employing sLSTM over mLSTM
Beck et al. [2024] present a comprehensive case for xLSTM blocks over earlier recurrent models. Here, we explain why, within those, we choose scalar-memory sLSTM blocks as our mixing primitive rather than mLSTM. The primary motivation is the state tracking via memory mixing. sLSTM preserves hidden-to-hidden recurrences in its gates and cell update, enabling conditional, history-aware updates. mLSTM removes these paths to allow full parallelization, which limits this state tracking capability.
Mechanisms of sLSTM and mLSTM.   The sLSTM block maintains the LSTM-style memory mixing, augmented with exponential gates and a stabilization term:
$c_{t}=f_{t}c_{t-1}+i_{t}z_{t},\qquad\qquad h_{t}=o_{t}g(c_{t}).$ (9)
Here, $f_{t},i_{t},o_{t}$ each depend on $h_{t-1}$ through recurrent matrices (Beck et al. [2024], Greff et al. [2017]), allowing the block to blend past and present conditionally. By contrast, mLSTM writes key-value pairs into a matrix memory:
$C_{t}=f_{t}C_{t-1}+i_{t}v_{t}k_{t}^{\top},$ (10)
with gates that depend only on the current input (no hidden-to-hidden paths) [Beck et al., 2024, Sec. 2.3]. This design eliminates memory mixing. Note that we use the scalar formulation for simplicity, whereas Sec. 2.1 directly presented the full vector-valued xLSTMs.
Importance of Latent Phases.   In long-horizon forecasting, tracking latent phases (e.g., heating $\leftrightarrow$ cooling; up $\leftrightarrow$ down) is critical. With sLSTM, gates can be trained such that, upon detection of a switch cue, the cell overwrites ($i_{t}\approx 1,f_{t}\approx 0$); otherwise, it retains ($i_{t}\approx 0,f_{t}\approx 1$). Because the gates read $h_{t-1}$, the decision is conditional on the current internal state. In mLSTM, gates cannot use $h_{t-1}$, meaning updates become fixed linear functionals of the input history. Empirically, architectures lacking memory mixing (mLSTM, Mamba, Transformers) fail simple state-tracking tasks such as parity, as shown by Merrill et al. [2024] and Beck et al. [2024].
Design Consequences.   For long-horizon forecasting, state-tracking capacity outweighs raw matrix-memory capacity, precluding the direct learning of inter-token relations during joint mixing. We, thus, exclusively employ sLSTM in xLSTM-Mixer. In our ablations, replacing sLSTM with mLSTM, GRU, or LSTM consistently harms long-horizon accuracy, particularly at 336/720 steps (cf. Tab. 3). Prior and concurrent work supports this finding: xLSTMTime uses sLSTM on smaller forecasting datasets where precise phase tracking dominates and opts for mLSTM only when raw storage capacity is paramount [Alharthi and Mahmood, 2024]. TiRex likewise attributes gains to retaining state tracking [Auer et al., 2025].
Appendix C Implementation Details
Experimental Details.
@@ Appendix C Implementation Details
Our codebase is implemented in Python 3.11, leveraging PyTorch version 2. [Paszke et al., 2019] in combination with Lightning version 2.4222https://lightning.ai/pytorch-lightning for model training and optimization. We used the custom CUDA implementation [^3] https://github.com/NX-AI/xlstm for sLSTM, which relies on NVIDIA Compute Capability $\geq$ 8.0. Thus, our experiments were conducted on a single NVIDIA A100 80GB GPU. The majority of our baseline implementations, along with data loading and preprocessing steps, are adapted from the Time-Series-Library [^4] https://github.com/thuml/Time-Series-Library of Wang et al. [2024b]. For GIFT-Eval [Aksu et al., 2024], we integrated xLSTM-Mixer into the official evaluation harness via a GluonTS-style [Alexandrov et al., 2020] estimator. For xLSTMTime, we used code based on the official repository [^5] https://github.com/muslehal/xLSTMTime. We employ Captum [^6] https://captum.ai  [Kokhlikyan et al., 2020] to compute the SHAP values used in model analysis. We used scikit-posthocs [Terpilowski, 2019] for significance analyses.
Training and Hyperparameters.
@@
We optimized xLSTM-Mixer in 32 bits for up to 60 epochs with a cosine-annealing scheduler with the Adam optimizer [Kingma and Ba, 2015], using $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and no weight decay. Hyperparameter (HP) tuning was conducted using Optuna [Akiba et al., 2019] with the choices provided in Tab. 4. We optimized for the L1 forecast error, also known as the Mean Absolute Error (MAE). To further stabilize the training process, gradient clipping with a maximum norm of $1.0$ was applied. All experiments were run with the three different random seeds $\{2021,2022,2023\}$. For most models, the initial publications already provide HPs appropriate for the well-known datasets, where we thus directly adopt these results. For xLSTMTime, this was not the case, and we were not able to fully reproduce the results in the paper despite our best efforts (cf. Sec. 5). We still present the better results from Alharthi and Mahmood [2024] so as not to erroneously underestimate the method. For the datasets ETTh2, ETTm1, and ETTm2 without such results, we ran individual HP searches analogously to our method to ensure a fair comparison.
@@
Forecasting Metrics.
We follow common practice in the literature [Wu et al., 2021, Wang et al., 2024a] for maximum comparability and, therefore, evaluate deterministic long-term forecasting of all models on the mean absolute error (MAE), mean squared error (MSE). For probabilistic evaluation (as in GIFT-Eval [Aksu et al., 2024]), we use the Mean Absolute Scaled Error (MASE) and the Continuous Ranked Probability Score (CRPS). All metrics are computed per variate and averaged over all variates. Formally,
[…]
The CRPS for predictive CDF $\hat{F}$ and outcome $y$ is
$\operatorname{CRPS}(\hat{F},y)$ $=\int_{-\infty}^{\infty}\!\big(\hat{F}(z)-\mathbbm{1}\{y\leq z\}\big)^{2}\,\mathrm{d}z\;=\;\mathbb{E}\lvert X-y\rvert\;-\;\tfrac{1}{2}\mathbb{E}\lvert X-X^{\prime}\rvert$,
where $\bm{y}$ are the targets, $\hat{\bm{y}}$ the predictions; $H$ is the horizon length, $T$ the in-sample length, $m$ a seasonal period, $d$ the MASE scaling constant (computed per variate), $\hat{F}$ the predictive CDF, and $X,X^{\prime}\!\sim\!\hat{F}$ i.i.d. When forecasts are provided as quantiles $\{\hat{q}_{t,\tau_{k}}\}_{k=1}^{K}$, we use the standard discrete approximation to CRPS via the pinball loss:
$\widehat{\operatorname{CRPS}}_{t}$ $=\frac{2}{K}\sum_{k=1}^{K}\rho_{\tau_{k}}\!\big(y_{t}-\hat{q}_{t,\tau_{k}}\big),\qquad\rho_{\tau}(u)=(\tau-\mathbb{1}\{u<0\})\,u,$
$\overline{\operatorname{CRPS}}$ $=\frac{1}{H}\sum_{t=1}^{H}\widehat{\operatorname{CRPS}}_{t}.$
Appendix D Benchmark Datasets
Tab. 5 provides an overview of the datasets we used to compare xLSTM-Mixer with other time series forecasting models. The last column shows the range of Hurst exponents [Hurst, 1951] over the variates measuring long-term patterns. The larger the values are over 0.5, the more long-term patterns are in the time series.
Dataset Source Domain Horizons Sampling #Variates Hurst exp.
Weather Zhou et al. [2021] Weather 96–720 10 min 21 0.333–1.000
Electricity Zhou et al. [2021] Power Usage 96–720 1 hour 321 0.555–1.000
Traffic Wu et al. [2021] Traffic Load 96–720 1 hour 862 0.162–1.000
ETT Zhou et al. [2021] Power Production 96–720 15&60 min 7 0.906–1.000
Appendix E Full Results for Long-Term Forecasting
Table 6: Full long-term forecasting results for Tab. 1. Avg is averaged from all four prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted bold red and second-best blue and underlined. Wins for each model out of all 28 settings are shown at the bottom.
Models Recurrent Mixer MLP State Space Transformer Convolutional Pretrained*
xLSTM-
@@ Appendix E Full Results for Long-Term Forecasting
TimeMix.++
TimeMix.